OpenStudy (anonymous):

is the square of the integral of a function same as the integral of function square?

OpenStudy (anonymous):

If I understand correctly : \[\left[ \int\limits\limits_{}^{}x dx \right]^2 \neq\int\limits\limits_{}^{}x^2 dx\]

OpenStudy (anonymous):

the same can be said about sqrtrt right?